OpenStudy (anonymous):

Alex designed the model of a cylindrical water tank to be installed in a park. The model has a radius of 4 cm and a height of 12 cm. The dimensions of the actual water tanks that will be installed in the park are 5 times larger than the model. The total surface area of the actual water tanks, rounded to the nearest whole number, is _________ cm2.

OpenStudy (anonymous):

can i get steps on how you solve it please

OpenStudy (shane_b):

First, do you know how to get the surface area of a cylinder?

OpenStudy (anonymous):

no i do not

OpenStudy (shane_b):

Ok, that's just a formula: \[SA_{cylinder}=top~area+bottom~area+side~area\]so:[\[SA_{cylinder}=2(area~of~top) + 2(perimeter~of~top)*height\]

OpenStudy (shane_b):

Finding the area of the top and bottom is easy: \[A_{top~or~bottom}=\pi r^2\]The area of the side will be the circumference times the height:\[A_{side}=(2 \pi r)(height)\]

OpenStudy (shane_b):

Put that all together to get: \[SA_{cylinder}=2\pi r^2 + 2\pi rh\]

OpenStudy (shane_b):

Are you with me so far?

OpenStudy (anonymous):

sorta kinda....

OpenStudy (shane_b):

What part is confusing you at the moment? We'll get to the calculation next but that's the easy part.

OpenStudy (anonymous):

o' ok i got it now.....i just had to re-read. i was looking at the surface area formula and what u wrote then saw the area formual and was like huh? then i got farther down and saw that you put it together so surface formula and area formula together which is how u got S.A.= 2pir2 + 2pirh

OpenStudy (shane_b):

Ok, great. Using the last equation above, calculate the SA of the model cylinder and tell me what you get please.

OpenStudy (anonymous):

ok hang on a sec.

OpenStudy (anonymous):

2(3.14)4cubed is....100.48 2(3.14)4x12 is 301.44 which equals 401.92?

OpenStudy (shane_b):

Here:\[SA=2\pi(4^2)+2\pi(4)(12) \approx402\]

OpenStudy (shane_b):

That's in square inches of course.

OpenStudy (shane_b):

If you're good with that, let's do the last part.

OpenStudy (anonymous):

was my math correct though?

OpenStudy (shane_b):

Your math is correct but in the first line you through me for a loop because you say cubed and not squared :)

OpenStudy (anonymous):

o sorry lol :)

OpenStudy (shane_b):

So now you know the formula and how to apply it. The last part of the question states that the dimensions of the actual tanks are 5 times larger. So what do you think you need to do?

OpenStudy (anonymous):

do i take 401.92 and multiply it by 5?

OpenStudy (shane_b):

No..that would be a silly mistake. All you need to do is multiply the dimensions by 5 and THEN apply the formula again.

OpenStudy (anonymous):

ok so i multiply 100.48 by 5 and 301.44 by 5?

OpenStudy (shane_b):

No...that's equivalent to what I said not to do :). Multiply the dimensions given in the problem by 5...and then do the formula.

OpenStudy (anonymous):

so if im understanding corrrectly...im going to multiply 4(radius) by 5 the (height) 12 by 5 and then put those in the formula like before

OpenStudy (shane_b):

Exactly

OpenStudy (anonymous):

oki doki hang on

OpenStudy (anonymous):

2512 946521.6 which equals 948133.6?

OpenStudy (shane_b):

I got 10053in^2 (rounded to the nearest whole number). Try it again...and I'll check my answer.

OpenStudy (shane_b):

I got 10048 this time but only because I used 3.14 instead of the longer version of pi.

OpenStudy (shane_b):

\[2(3.14)20^2+2(3.14)(20)(60)=10048\]

OpenStudy (anonymous):

ok this is what im doing... 2(3.14)20 squared=2512 2(3.14) 20 times 60 = 7536 which equals 10048

OpenStudy (anonymous):

ok i see what i did....i used the 20squared in the second half of the equation lol

OpenStudy (anonymous):

insead of just 20

OpenStudy (shane_b):

Glad you found the mistake! Ok, so that's the final answer. Note that if you had just multiple 402x5 you would have came up with 2010...which is far different than 10,000. The reason you can't just multiply your first answer time 5 is because there are exponents in the formula. If those weren't there, it would have worked.

OpenStudy (shane_b):

You didn't even need to do the first answer...but we did it so you could understand the formula.

OpenStudy (anonymous):

ohhh ok well i totaly understand it now...thank you sooo much you helped me alot :)

OpenStudy (shane_b):

cool, good luck :)

OpenStudy (anonymous):

hang on wait....if it says use 22/7 for pi would it still be the same? or how would i do that

OpenStudy (anonymous):

@Shane_B if i had to use 22/7 for pi instead of 3.14 would i end with 10057.142857?

OpenStudy (shane_b):

That's correct

OpenStudy (anonymous):

ok cool thanks (: